How many liters of a 12% acid solution and a 20% acid solution should be mixed in order to obtain 4 liters of a 15% acid solution?
Answer by mananth(16949)

percent ---------------- quantity
Acid solution I 12 ---------------- x liters
Acid solution II 20 ------ 4 - x liters
Mixture 15 ---------------- 4

12x+20(4-x)=15*4
1200x+2000(4-x) =6000
1200x+8000-2000 x=6000
1200x- 2000x=6000 -8000
-800 x = -2000
/ -800
x=2.5 liters 12.00% Acid solution I
1.5 liters 20.00% Acid solution II

m.ananth@hotmail.ca